OpenStudy (anonymous):

Use dimensional analysis to show that E has units of energy (Joules) in E= h (c/lambda)

OpenStudy (rogue):

\[E = \frac {hc}{\lambda}\]The unit for Planck's constant is J*s. The speed of light is measured in meters per seconds (m/s). Wavelength is in meters. \[E = \frac {J * s \times m/s}{m} = J\]